# Comparison of three methods for triggering core-collapse supernova explosions in spherical symmetry

### v1

@@ 1 Introduction
While the true nature of CCSNe involves multi-dimensional hydrodynamic instabilities such as convection and energy-dependent three-flavor neutrino transport, spherical symmetry (1D) and various kinds of approximations for the triggering mechanism of the explosions, often sacrificing the treatment of neutrinos and the neutron-star physics, serve as a useful starting point. These simplifications permit time-dependent simulations in one dimension, i.e., just with a radius dependence, with much less computational demands. Therefore they have been widely employed, in particular, for determining CCSN nucleosynthesis (for an overview, see, e.g., Umeda & Yoshida, 2017, and references therein) and CCSN light curves and spectra (e.g., Sukhbold et al., 2016; Goldberg et al., 2019; Goldberg & Bildsten, 2020; Teffs et al., 2020a, b; Dessart et al., 2021a, b; Curtis et al., 2021; Dessart et al., 2023). Beyond such applications, approximate treatments also allow for systematic studies of large sets of CCSN calculations that are essential for a better understanding of the fundamental dependencies of stellar core collapse on crucial nuclear and particle physics, thus aiding the development of more comprehensive multi-dimensional models. Moreover, 1D simplifications also facilitate the advancement of our knowledge of stellar evolution, chemical element formation, and of the chemical enrichment history of galaxies (e.g., Timmes et al., 1995; Matteucci, 2003; Hayden et al., 2015; Kobayashi et al., 2020; Wirth et al., 2021). Similarly, deeper insights into the statistics and birth-mass distributions of resulting compact objects, i.e., neutron stars and black holes, gain from such studies (e.g., Zhang et al., 2008; O’Connor & Ott, 2011; Ugliano et al., 2012; Pejcha & Thompson, 2015; Ertl et al., 2016; Sukhbold et al., 2016; Ebinger et al., 2019, 2020; Ertl et al., 2020; Woosley et al., 2020; Boccioli & Fragione, 2024). Thus they provide crucial information to refine models and simulations, contribute to our understanding of the Universe’s chemical evolution, and pave the way for multi-messenger astronomy including gravitational-wave and neutrino measurements that offer a more complete picture of these extraordinary astrophysical events.
@@
The thermal-bomb method was introduced by K. Nomoto and collaborators (e.g., Shigeyama et al., 1988; Hashimoto et al., 1989; Thielemann et al., 1990; Nakamura et al., 2001; Nomoto et al., 2006). It was also used in more recent works (e.g., Moriya et al., 2010; Sawada & Maeda, 2019), and it is the explosion trigger implemented in the 1D-codes MESA (Paxton et al., 2015) and SNEC (Morozova et al., 2015). The thermal bomb “mechanism” is also used for simulating hypernovae and low-energy explosions (e.g., Nomoto et al., 2006; Umeda & Nomoto, 2008; Nomoto et al., 2009). In hydrodynamic simulations, the thermal bomb can be implemented as a parametrized energy deposition term in the energy equation. By introducing this artificial release of thermal energy in a chosen volume (mass for a Lagrangian, and spatial domain for an Eulerian approach) and over a defined period of time, the aim, in the best case, is to more or less closely mimic the energy input expected by the true physical processes in the core of the CCSN. In our previous paper (Imasheva et al., 2023) we investigated different setups for such thermal bombs and came up with some general recommendations for suitable treatments and parameter choices, which will be mentioned in Section 2.2.3.
@@
Thermal bombs as well as pistons do not include neutrinos in their descriptions, which has consequences for the dynamics and the nucleosynthesis conditions in the CCSN models. In view of these limitations, the more sophisticated 1D neutrino engines to drive explosions by neutrino-energy deposition behind the CCSN shock in 1D hydrodynamic simulations were more recently introduced. These engines exist in various versions, which use different ingredients to achieve successful blast waves: enhanced postshock heating by electron neutrinos and antineutrinos ($\nu_{e}$ and $\bar{\nu}_{e}$) via a parametric increase of their absorption rates on neutrons and protons (O’Connor & Ott, 2011); a simple spherical and grey approximation of neutrino transport and increased $\nu_{e}$ and $\bar{\nu}_{e}$ luminosities from the proto-neutron star’s high-density core (P-HOTB; Ugliano et al., 2012; Ertl et al., 2016; Sukhbold et al., 2016); a neutrino-driven wind, which is injected at the inner grid boundary and blows out the overlying layers of the progenitor (Pejcha & Thompson, 2015); energy-dependent, three-flavor, general-relativistic (GR) Boltzmann neutrino transport with a parameterization of additional postshock heating linked to the heavy-lepton neutrino ($\nu_{\mu}$, $\nu_{\tau}$, $\bar{\nu}_{\mu}$, $\bar{\nu}_{\tau}$) luminosities (PUSH; Perego et al., 2015); or a 1D approximation of turbulent convection in the neutrino-heated postshock layer to aid explosions in 1D GR neutrino-hydrodynamic models (STIR; Couch et al., 2020; Boccioli et al., 2021). In our present work we obtain neutrino-driven explosions in 1D CCSN simulations by the P-HOTB engine.
@@ 3.2 Nickel production
The tracer material mainly forms when neutrino-heated, neutron-rich ejecta that are initially at NSE temperatures experience expansion cooling and thus nucleon recombination to $\alpha$-particles and heavy nuclei during the freeze-out or alpha-rich freeze-out from NSE. Therefore the production of the tracer depends on the electron fraction in this material, which, however, is uncertain in our P-HOTB models because of the approximate treatment of the neutrino transport and a highly sensitive dependence of $Y_{e}$ on the radiated neutrino luminosities and spectra. For this reason some of the tracer material could be 56Ni instead, which motivates us to consider the mass of 56Ni as a lower limit to the production of this isotope and the summed yields of 56Ni and tracer, $M_{\rm Ni+tr}$, as an upper limit for the ejected 56Ni mass (see the discussion by Sukhbold et al., 2016). We note in passing, however, that 1D neutrino-driven explosions are not able to adequately capture the nucleosynthesis conditions in the innermost ejecta of 3D models of neutrino-driven explosions (e.g., Wongwathanarat et al., 2013, 2017; Bruenn et al., 2016; Wanajo et al., 2018; Sieverding et al., 2023; Wang & Burrows, 2024a, b). Because of differences in the electron fraction, entropy, and expansion history, precise predictions of iron-group and trans-iron abundances cannot be expected from the 1D simulations, and even some of the matter in 1D surviving as $\alpha$-particles during the freeze-out from NSE might instead form 56Ni in 3D explosions (we refer to the corresponding discussion by Ertl et al., 2020). Therefore, at best, the overall magnitude of the yields and their general trends (e.g., variations with progenitor and explosion energy) estimated on grounds of 1D neutrino-driven CCSN simulations might be similar to the multi-dimensional results.
@@ 4 Explosion trends across progenitors
Moreover, for the neutrino-driven explosions (Figure 10, left panel), there is a clear correlation between $M_{\rm Ni+tracer}$ (cyan dashed line) and the explosion energy at 2 s (red line): the higher the energy is, the higher the final mass of nickel plus tracer is, both having peaks for progenitors of 21.0, 26.6, and 19.8 $M_{\odot}$. This correlation can be understood by the fact that the energy of neutrino-driven explosions is provided by energy deposition of neutrinos in a part of the ejecta that in the subsequent freeze-out becomes iron-group (and trans-iron) elements, including nickel and tracer material, and, possibly, a remaining mass fraction of $\alpha$-particles (helium).

### v2

@@ 1 Introduction
While the true nature of CCSNe involves multi-dimensional hydrodynamic instabilities such as convection and energy-dependent three-flavor neutrino transport, spherical symmetry (1D) and various kinds of approximations for the triggering mechanism of the explosions, often sacrificing the treatment of neutrinos and the neutron-star physics, serve as a useful starting point. These simplifications permit time-dependent simulations in one dimension, i.e., just with a radius dependence, with much less computational demands. Therefore they have been widely employed, in particular, for determining CCSN nucleosynthesis (for an overview, see, e.g., Umeda & Yoshida, 2017, and references therein) and CCSN light curves and spectra (e.g., Sukhbold et al., 2016; Goldberg et al., 2019; Goldberg & Bildsten, 2020; Limongi & Chieffi, 2020; Teffs et al., 2020a, b; Dessart et al., 2021a, b; Curtis et al., 2021; Dessart et al., 2023). Beyond such applications, approximate treatments also allow for systematic studies of large sets of CCSN calculations that are essential for a better understanding of the fundamental dependencies of stellar core collapse on crucial nuclear and particle physics, thus aiding the development of more comprehensive multi-dimensional models. Moreover, 1D simplifications also facilitate the advancement of our knowledge of stellar evolution, chemical element formation, and of the chemical enrichment history of galaxies (e.g., Timmes et al., 1995; Matteucci, 2003; Hayden et al., 2015; Kobayashi et al., 2020; Wirth et al., 2021; Roberti et al., 2024; Jost et al., 2025). Similarly, deeper insights into the statistics and birth-mass distributions of resulting compact objects, i.e., neutron stars and black holes, gain from such studies (e.g., Zhang et al., 2008; O’Connor & Ott, 2011; Ugliano et al., 2012; Pejcha & Thompson, 2015; Ertl et al., 2016; Sukhbold et al., 2016; Ebinger et al., 2019, 2020; Ertl et al., 2020; Woosley et al., 2020; Boccioli & Fragione, 2024; Ugolini et al., 2025). Thus they provide crucial information to refine models and simulations, contribute to our understanding of the Universe’s chemical evolution, and pave the way for multi-messenger astronomy including gravitational-wave and neutrino measurements that offer a more complete picture of these extraordinary astrophysical events.
@@
The thermal-bomb method was introduced by K. Nomoto and collaborators (e.g., Shigeyama et al., 1988; Hashimoto et al., 1989; Thielemann et al., 1990; Nakamura et al., 2001; Nomoto et al., 2006). It was also used in more recent works (e.g., Moriya et al., 2010; Sawada & Maeda, 2019; Ko et al., 2022), and it is the explosion trigger implemented in the 1D-codes MESA (Paxton et al., 2015) and SNEC (Morozova et al., 2015). The thermal bomb “mechanism” is also used for simulating hypernovae and low-energy explosions (e.g., Nomoto et al., 2006; Umeda & Nomoto, 2008; Nomoto et al., 2009). In hydrodynamic simulations, the thermal bomb can be implemented as a parametrized energy deposition term in the energy equation. By introducing this artificial release of thermal energy in a chosen volume (mass for a Lagrangian, and spatial domain for an Eulerian approach) and over a defined period of time, the aim, in the best case, is to more or less closely mimic the energy input expected by the true physical processes in the core of the CCSN. In our previous paper (Imasheva et al., 2023) we investigated different setups for such thermal bombs and came up with some general recommendations for suitable treatments and parameter choices, which will be mentioned in Section 2.2.3.
@@
Thermal bombs as well as pistons do not include neutrinos in their descriptions, which has consequences for the dynamics and the nucleosynthesis conditions in the CCSN models. In view of these limitations, the more sophisticated 1D neutrino engines to drive explosions by neutrino-energy deposition behind the CCSN shock in 1D hydrodynamic simulations were more recently introduced. These engines exist in various versions, which use different ingredients to achieve successful blast waves: enhanced postshock heating by electron neutrinos and antineutrinos ($\nu_{e}$ and $\bar{\nu}_{e}$) via a parametric increase of their absorption rates on neutrons and protons (O’Connor & Ott, 2011; Jost et al., 2025); a simple spherical and grey approximation of neutrino transport and increased $\nu_{e}$ and $\bar{\nu}_{e}$ luminosities from the proto-neutron star’s high-density core (P-HOTB; Ugliano et al., 2012; Ertl et al., 2016; Sukhbold et al., 2016); a neutrino-driven wind, which is injected at the inner grid boundary and blows out the overlying layers of the progenitor (Pejcha & Thompson, 2015); energy-dependent, three-flavor, general-relativistic (GR) Boltzmann neutrino transport with a parameterization of additional postshock heating linked to the heavy-lepton neutrino ($\nu_{\mu}$, $\nu_{\tau}$, $\bar{\nu}_{\mu}$, $\bar{\nu}_{\tau}$) luminosities (PUSH; Perego et al., 2015); or a 1D approximation of turbulent convection in the neutrino-heated postshock layer to aid explosions in 1D Newtonian and GR neutrino-hydrodynamic models (STIR; Couch et al., 2020; Boccioli et al., 2021; Barker et al., 2022). All of these neutrino-engine methods were also used in large sets of 1D explosion models. In our present work we obtain neutrino-driven explosions in 1D CCSN simulations by the P-HOTB engine.
@@ 3.2 Nickel production
The tracer material mainly forms when neutrino-heated, neutron-rich ejecta that are initially at NSE temperatures experience expansion cooling and thus nucleon recombination to $\alpha$-particles and heavy nuclei during the freeze-out or alpha-rich freeze-out from NSE. Therefore the production of the tracer depends on the electron fraction in this material, which, however, is uncertain in our P-HOTB models because of the approximate treatment of the neutrino transport and a highly sensitive dependence of $Y_{e}$ on the radiated neutrino luminosities and spectra. For this reason some of the tracer material could be 56Ni instead, which motivates us to consider the mass of 56Ni as a lower limit to the production of this isotope and the summed yields of 56Ni and tracer, $M_{\rm Ni+tr}$, as an upper limit for the ejected 56Ni mass (see the discussion by Sukhbold et al., 2016). We note in passing, however, that 1D neutrino-driven explosions are not able to adequately capture the nucleosynthesis conditions in the innermost ejecta of 3D models of neutrino-driven explosions (e.g., Wongwathanarat et al., 2013, 2017; Bruenn et al., 2016; Wanajo et al., 2018; Sieverding et al., 2023; Wang & Burrows, 2024a, b). Because of differences in the electron fraction, entropy, and expansion history, precise predictions of iron-group and trans-iron abundances cannot be expected from the 1D simulations. For example, mass elements ejected at a certain time in 1D models possess a single value of $Y_{e}$, entropy, and expansion time scale, whereas in 3D the ejected matter has a distribution of different conditions at each moment. Moreover, even some of the matter surviving as $\alpha$-particles in 1D during the freeze-out from NSE might instead form 56Ni in 3D explosions, because the entropies in neutrino-heated ejecta are overestimated in 1D models (we refer to the corresponding discussion by Ertl et al., 2020). Therefore, at best, the overall magnitude of the yields and their general trends (e.g., variations with progenitor and explosion energy) estimated on grounds of 1D neutrino-driven CCSN simulations might be similar to the multi-dimensional results.
@@ 4 Explosion trends across progenitors
Moreover, for the neutrino-driven explosions (Figure 10, left panel), there is a clear correlation between $M_{\rm Ni+tracer}$ (cyan dashed line) and the explosion energy at 2 s (red line): the higher the energy is, the higher the final mass of nickel plus tracer is, both having peaks for progenitors of 21.0, 26.6, and 19.8 $M_{\odot}$. This correlation can be understood by the fact that the energy of neutrino-driven explosions is provided by energy deposition of neutrinos in a part of the ejecta that in the subsequent freeze-out becomes iron-group (and trans-iron) elements, including nickel and tracer material, and, possibly, a remaining mass fraction of $\alpha$-particles (helium). We repeat that some of the tracer material could well be 56Ni and possibly even a part of the $\alpha$-particles not recombining in the neutrino-heated ejecta of 1D models could well form 56Ni in 3D explosions of similar energies (see the discussion in Sect. 3.2 and references given there). [^1] In this context we point to 3D CCSN results by Burrows et al. (2024). In their Fig. 13, left panel, the 56Ni masses correlate with explosion energies. In the energy range from 0.5 B to 2.0 B, which is relevant for our study, this correlation displays considerable scatter with roughly the same 56Ni mass occurring for different explosion energies. The 56Ni masses are between $\sim$0.06 $M_{\odot}$ and $\sim$0.17 $M_{\odot}$. This is not too different from the 56Ni plus tracer masses of our neutrino-engine models, which range between $\sim$0.05 $M_{\odot}$ and $\sim$0.13 $M_{\odot}$ with a correlation that also shows some scatter, i.e., similar 56Ni masses are possible for different explosion energies (see Figure 10, left panel).